OpenStudy (anonymous):

Evaluate AB/CD for A = 5, B = -2, C = 4 and D = -6.

OpenStudy (anonymous):

To clarify: A * B ---- C * D ?

OpenStudy (anonymous):

5 * -2 / 4 * -6

OpenStudy (anonymous):

looks like you can cancel some stuff

OpenStudy (anonymous):

\[\frac{AB}{CD } =\frac{A \times B}{C \times D } =\frac{5 \times (-2)}{4\times (-6) } \] \[=\frac{5 }{4\times 3} =\frac{5 }{12} \]

OpenStudy (anonymous):

hey , i thought you can't give direct answers